OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! Which is the simplified form of 18q^4r/6q^3r^5

OpenStudy (anonymous):

ANSWER: 3q/r^4. 18/6 = 3 q^4/q^3 = q r/r^5= 1/r^4 You just divide each part of the of the numerator by the denominator.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

You're Welcome, do you understand how to do it now?

OpenStudy (anonymous):

Yes I do!

OpenStudy (anonymous):

@wmckinely Could you help me with another?

OpenStudy (anonymous):

Well post it, and I'll see if I can:)